# Burer-Monteiro factorizability of nuclear norm regularized optimization

### v1

@@ 4 Approximate stationary points of $f$
$\begin{split}0&\leq-2d_{r+1}+2\lambda+\nabla^{2}h(\bar{U}\bar{V}^{\top})\left[\sigma_{r}(\bar{U})R\begin{bmatrix}0_{r\times r}&e_{r}e_{1}^{\top}\\ e_{1}e_{r}^{\top}&0_{(m-r)\times(n-r)}\end{bmatrix}P^{\top}\right]^{2}\\ &\leq-2d_{r+1}+2\lambda+2\sigma_{r}^{2}(\bar{U})L=-2d_{r+1}+2\lambda+2\sigma_{r}(\bar{U}\bar{V}^{\top})L,\end{split}$
where the second inequality follows from the fact that $\left\|\sigma_{r}(\bar{U})R\begin{bmatrix}0_{r\times r}&e_{r}e_{1}^{\top}\\ e_{1}e_{r}^{\top}&0_{(m-r)\times(n-r)}\end{bmatrix}P^{\top}\right\|_{F}=\sqrt{2}\sigma_{r}(\bar{U})$ and the definition of $L$, and the equality follows from Remark 3.3(ii). Hence, $d_{r+1}\leq\lambda+L\,\sigma_{r}(\bar{U}\bar{V}^{\top})$. In addition, we can further compute that
$\,\inf_{Y\in\partial f(\bar{U}\bar{V}^{\top})}\|Y\|_{2}=\inf_{S\in\partial\|\bar{U}\bar{V}^{\top}\|_{*}}\|\nabla h(\bar{U}\bar{V}^{\top})+\lambda S\|_{2}$
$=$ $\,\inf_{\|W\|_{2}\leq 1}\left\|-R\begin{bmatrix}\lambda I_{r}&0_{r\times(n-r)}\\ 0_{(m-r)\times r}&\widetilde{\rm Diag}(d_{r+1},\dots,d_{m})\end{bmatrix}P^{\top}+\lambda R\begin{bmatrix}I_{r}&0_{r\times(n-r)}\\ 0_{(m-r)\times r}&W\end{bmatrix}P^{\top}\right\|_{2}$
$=$ $\,\inf_{\|W\|_{2}\leq\lambda}\|\widetilde{\rm Diag}(d_{r+1},\dots,d_{m})-W\|_{2}=\max\{d_{r+1}-\lambda,0\}\leq L\sigma_{r}(\bar{U}\bar{V}^{\top}),$
where the first equality follows from the definition of $f$, the second follows from 2.1 and 3.2, the third holds upon making a simple change of variables, the fourth makes use of the unitary invariance of $\|\cdot\|_{2}$, [23, Proposition 2.1] and the fact that $d_{r+1}\geq\cdots\geq d_{m}\geq 0$, and the inequality holds because $d_{r+1}\leq\lambda+L\,\sigma_{r}(\bar{U}\bar{V}^{\top})$. This completes the proof. ∎
In view of 4.1, for a second-order stationary point $(\bar{U},\bar{V})$ of $F_{r}$ in 1.5, by setting $\bar{X}=\bar{U}\bar{V}^{\top}$, we see that the smaller $\sigma_{r}(\bar{X})$ is, the closer $\bar{X}$ is to being a stationary point of $f$ in 1.4. When $\sigma_{r}(\bar{X})=0$, we see that $\bar{X}$ is a stationary point of $f$, as proved in [2]; see, also [41, Section 3], and [14, 3, 4] for similar results.
Corollary 4.2.
Our next goal is to characterize rank-deficient second-order stationary points of $F_{r}$ in (1.5) under the convexity of the function $h$. To do so, we need the following lemma.
Lemma 4.3.
Let $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ satisfy $\bar{U}^{\top}\bar{U}=\bar{V}^{\top}\bar{V}$ and that $\bar{U}\bar{V}^{\top}$ is a local (global) minimizer of problem (1.4). Then, $(\bar{U},\bar{V})$ is a local (global) minimizer of problem 1.5.
Proof.
We only prove the statement for local minimizers, as the proof of the statement for global minimizers is the same. Let $(P,Q)\in\mathcal{O}_{\bar{V}}$. By Lemma 3.1, there exists $R\in\mathcal{O}^{m}$ such that $(R,Q)\in\mathcal{O}_{\bar{U}}$. Lemma 3.1 also asserts that $\sigma(\bar{V})=\sigma(\bar{U})$. Therefore,
$\|\bar{U}\bar{V}^{\top}\|_{*}=\|R\widetilde{\rm Diag}(\sigma(\bar{U}))Q^{\top}Q\widetilde{\rm Diag}(\sigma(\bar{V}))P^{\top}\|_{*}=\sigma(\bar{U})^{\top}\sigma(\bar{V})=\|\sigma(\bar{U})\|_{2}^{2}=\frac{1}{2}(\|\bar{U}\|_{F}^{2}+\|\bar{V}\|_{F}^{2}),$
which implies that $F_{r}(\bar{U},\bar{V})=f(\bar{U}\bar{V}^{\top})$. Next, since $\bar{U}\bar{V}^{\top}$ is a local minimizer of $f$, by definition, there exists an $\epsilon>0$ such $f(X)\geq f(\bar{U}\bar{V}^{\top})$ for any $X$ satisfying $\|X-\bar{U}\bar{V}^{\top}\|\leq\epsilon$. By the continuity of the mapping $(U,V)\mapsto UV^{\top}$, there exists an $\epsilon^{\prime}>0$ such that $\|UV^{\top}-\bar{U}\bar{V}^{\top}\|_{F}\leq\epsilon$ whenever $\|(U,V)-(\bar{U},\bar{V})\|_{F}\leq\epsilon^{\prime}$. Consider any $(U,V)$ satisfying $\|(U,V)-(\bar{U},\bar{V})\|_{F}\leq\epsilon^{\prime}$. Then,
$\,F_{r}(U,V)=h(UV^{\top})+\frac{\lambda}{2}(\|U\|_{F}^{2}+\|V\|_{F}^{2})\geq h(UV^{\top})+\lambda\min_{U^{\prime}{V^{\prime}}^{\top}=UV^{\top}}\frac{\|U^{\prime}\|_{F}^{2}+\|V^{\prime}\|_{F}^{2}}{2}$
$=$ $\,h(UV^{\top})+\lambda\|UV^{\top}\|_{*}=f(UV^{\top})\geq f(\bar{U}\bar{V}^{\top})=F_{r}(\bar{U},\bar{V}),$
where the second equality follows from [34, Lemma 1]. Hence, $(\bar{U},\bar{V})$ is a local minimizer of $F_{r}$. This completes the proof. ∎
The theorem below shows that when the function $h$ in problem (1.4) is convex, a rank-deficient second-order stationary point of problem (1.5) does not only induce a stationary point for problem (1.4), but is also a global minimizer of problem (1.5).
Theorem 4.4 (Second-order stationary point with rank deficiency).
Consider problems (1.4) and (1.5) with convex $h$. For any $(U,V)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ satisfying $\mathrm{rank}(U)<r$, the following statements are equivalent:
  1. (i)
    $(U,V)$ is a second-order stationary point of problem (1.5);
  2. (ii)
    $UV^{\top}$ is a global minimizer of problem (1.4) and $U^{\top}U=V^{\top}V$;
  3. (iii)
    $(U,V)$ is a global minimizer of problem (1.5).
Proof.
It is trivial that (iii) implies (i). We then prove that (i) implies (ii). By 4.2, $UV^{\top}$ is a stationary point of problem (1.4). Since $h$ is convex, the objective function $f$ of problem (1.4) is also convex. The global minimality then follows from the stationarity of $UV^{\top}$, which together with [19, Proposition 2] implies that $U^{\top}U=V^{\top}V$. Finally, it follows directly from Lemma 4.3 that (ii) implies (iii). The proof is thus completed. ∎
5 Characterization of $r$-factorizability
In view of 4.4, any rank-deficient second-order stationary point of $F_{r}$ in 1.5 is a global minimizer of $f$ (1.4) when $h$ is convex. Consequently, if $h$ is convex and all second-order stationary points of (1.5) are rank-deficient, then $f$ is $r$-factorizable. Moreover, the discussion preceding 4.1 suggests that the rank conditions cannot be dropped in general.
This section approaches the problem from a different perspective. In particular, we aim to derive characterizations of $r$-factorizability. Our characterizations are based on a set of carefully chosen parameters, as described in the following definition.
Definition 5.1.
Let $L\in(0,\infty),~{}M\in(0,\infty]$, $\mu\geq 0$, $r^{*}\in[m]\cup\{0\}$, and $0\leq q\leq m-r^{*}$. We define $\mathfrak{S}(L,\mu,r^{*},M,q)$ to be the set of all $h\in C^{2}(\mathbb{R}^{m\times n})$ satisfying the following conditions:
  1. (i)
    The function $h(\cdot)-\frac{\mu}{2}\|\cdot\|_{F}^{2}$ is convex, and $\nabla h$ is $L$-Lipschitz continuous.
  2. (ii)
    There exists a global minimizer $X^{*}\in\mathbb{R}^{m\times n}$ of $f$ in (1.4) satisfying $\mathrm{rank}(X^{*})=r^{*}$, $\|X^{*}\|_{2}\leq M$ and $\|\nabla h(X^{*})\|_{*}\leq\lambda(r^{*}+q)$.
In view of the first-order optimality condition of 1.4 and 2.1, one can observe that any Lipschitz differentiable convex $C^{2}$ function with (1.4) being solvable belongs to $\mathfrak{S}(L,\mu,r^{*},M,q)$ for some $L$, $\mu$, $r^{*}$, $M$ and $q$. We will study conditions on these parameters so that the corresponding $\mathfrak{S}(L,\mu,r^{*},M,q)$ consists solely of $h$ whose corresponding $f$ in 1.4 is $r$-factorizable. We will first consider in Section 5.1 strongly convex loss functions (i.e., $\mu>0$), where our characterizations of the parameters bear similarity to the recent work [44], which also considered strongly convex loss functions. The general case will be studied in Section 5.2.
5.1 Lipschitz differentiable strongly convex $C^{2}$ loss functions
We first characterize the $r$-factorizability of Lipschitz differentiable strongly convex $C^{2}$ functions. It turns out that under strong convexity, the characterization can be made independent of the parameters $M$ and $q$. We thus consider the class $\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$. One can observe that any Lipschitz differentiable strongly convex $C^{2}$ function belongs to $\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$ for some $L$, $\mu$ and $r^{*}$. Our main result is the following theorem, which characterizes when $\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$ consists solely of functions $h$ whose corresponding $f$ in 1.4 is $r$-factorizable.
Theorem 5.2.
Let $r^{*}\in[m]\cup\{0\}$, $r\in[m]$, $\infty>L\geq\mu>0$ and $\kappa=\frac{L}{\mu}\geq 1$. If $r^{*}$, $r$ and $\kappa$ satisfy any of the following conditions:
  1. (1)
    $r^{*}=0$ or $r=m$,
  2. (2)
    $r>r^{*}$ and $\kappa=3$,
  3. (3)
    $r\geq r^{*}$ and $\kappa<3$,
  4. (4)
    $r\geq r^{*}$, $\kappa>3$ and $\frac{r-r^{*}}{1-\frac{4}{(\kappa-1)^{2}}}>\min\{r,m-r^{*}\}$,
then for all $h\in\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$, the corresponding $f$ in 1.4 is $r$-factorizable. Otherwise, there exists a quadratic $h\in\mathfrak{S}(L,\mu,r^{*},\infty,m-r^{*})$ such that the corresponding $f$ in 1.4 is not $r$-factorizable.
Proof.
This follows from B.1, B.2 and B.4. ∎
Note that Item (3) in 5.2 also appeared in [44, Corollary 1.2] to ensure that, in 1.7, all second-order stationary points $U$ of $\widetilde{h}$ satisfy that $UU^{\top}$ globally minimizes $h$ over ${\mathbb{S}}_{+}^{n}$.
5.2 Lipschitz differentiable convex $C^{2}$ loss functions
In this subsection, we characterize $r$-factorizability of $f$ in 1.4 with a Lipschitz differentiable convex $C^{2}$ function $h$. In view of the first-order optimality condition of 1.4 and 2.1, one can observe that any such $h$ with (1.4) being solvable belongs to $\mathfrak{S}(L,0,r^{*},M,q)$ for some $L$, $r^{*}$, $M$ and $q$. The main result here is similar to 5.2. It characterizes when $\mathfrak{S}(L,0,r^{*},M,q)$ consists solely of functions $h$ whose corresponding $f$ in 1.4 is $r$-factorizable.
Theorem 5.3.
Assume $r^{*}\in[m]\cup\{0\}$ and $r\in[m]$. Let $q\in[0,m-r^{*}]$, $L$, $M\in(0,\infty)$ and $W^{*}$ be the optimal value of the optimization problem:
$\sup_{d\in\mathbb{N}_{0}}~{}(q-d)(1-\frac{q}{d})+\frac{(q-d+LM(r^{*}-r+d)/\lambda)^{2}}{4(r^{*}-r+d)}$ (5.1)
$s.t.~{}~{}~{}~{}\max\{r-r^{*}+1,1\}\leq d\leq\min\{r,m-r^{*}\},~{}q-d+LM(r^{*}-r+d)/\lambda>0.$
If $r$, $r^{*}$, $L$, $M$, $q$, $W^{*}$ satisfy any of the following conditions:
  1. (1)
    $r^{*}=0$,
  2. (2)
    $r^{*}>0$, $r\geq r^{*}+\lfloor q\rfloor$, and $W^{*}<0$,
then for all $h\in\mathfrak{S}(L,0,r^{*},M,q)$, the corresponding $f$ in 1.4 is $r$-factorizable. Otherwise, there exists a quadratic $h\in\mathfrak{S}(L,0,r^{*},M,q)$ such that the corresponding $f$ in 1.4 is not $r$-factorizable.
Remark 5.4.
In view of the inequality constraint involving $q$ in (5.1), one can see that the optimal value of 5.1 is decreasing in $\lambda$ and increasing in $M$. In addition, when $r\geq r^{*}+\lfloor q\rfloor$, the constraint in (5.1) implies that $d\geq r-r^{*}+1\geq\lfloor q\rfloor+1\geq q$. This further implies that the optimal value of 5.1 is increasing in $q\in[0,\min\{d,m-r^{*}\}]$. Therefore, the $W^{*}$ in Theorem 5.3 is more likely negative when $M$ and $q$ are small and $\lambda$ is large.
Proof for 5.3.
This follows from C.1, C.2 and Lemma C.4. ∎
Appendix A Pseudo-stationarity
Our main strategy for characterizing the $r$-factorizability is to analyze the pseudo-stationary points of $f$ in (1.4), which we define as follows.
Definition A.1 (Pseudo-stationarity).
A matrix $X\in\mathbb{R}^{m\times n}$ is said to be a pseudo-stationary point of $f$ in (1.4) if there exist $(R,P)\in\mathcal{O}_{X}$ and $d\in\mathbb{R}^{m}_{+}$ such that $-\nabla h(X)=R\widetilde{\rm Diag}(d)P^{\top}$ and $d_{1}=\cdots=d_{s}=\lambda$, where $s=\mathrm{rank}(X)$.
Definition A.1 is motivated by 2.1 and the first-order optimality condition of 1.4. Indeed, any pseudo-stationary point with the $d$ in Definition A.1 satisfying $\lambda\geq\max\{d_{s+1},\ldots,d_{m}\}$ is, in fact, a stationary point of $F_{r}$ in 1.5. Here, we present several lemmas concerning pseudo-stationary points that are useful for proving the main results in Section 5.
Lemma A.2.
Let $X_{1},X_{2}$ be two pseudo-stationary points of $f$ in 1.4 in the sense of Definition A.1, i.e., there exist $R_{1}$, $R_{2}\in\mathcal{O}^{m}$ and $P_{1}$, $P_{2}\in\mathcal{O}^{n}$ such that
$X_{1}=R_{1}\begin{bmatrix}\Sigma_{1}&0_{m\times(n-m)}\end{bmatrix}P_{1}^{\top},~{}-\nabla h(X_{1})=R_{1}\begin{bmatrix}D_{1}&0_{m\times(n-m)}\end{bmatrix}P_{1}^{\top},$ (A.1)
where $\Sigma_{i}=\mathrm{diag}(\sigma_{1}(X_{i}),\dots,\sigma_{m}(X_{i}))\in\mathbb{R}^{m\times m}$ and $D_{i}=\mathrm{diag}(d_{1}^{i},\dots,d_{m}^{i})\in\mathbb{R}^{m\times m}_{+}$ with $d^{i}_{1}=\dots=d^{i}_{\mathrm{rank}(X_{i})}=\lambda$ for $i=1,2$. Assume that $h$ in 1.4 satisfies that $h(\cdot)-\frac{\mu}{2}\|\cdot\|_{F}^{2}$ is convex for some $\mu\geq 0$, and $\nabla h$ is Lipschitz continuous with modulus $L\geq\mu$. Then, we have
$\max_{\tau\in\mathfrak{P}_{m}}$ $\sum_{i=1}^{m}(L\sigma_{i}(X_{1})+d^{1}_{i})(\mu\sigma_{\tau(i)}(X_{2})+d^{2}_{\tau(i)})+\sum_{i=1}^{m}(\mu\sigma_{i}(X_{1})+d^{1}_{i})(L\sigma_{\tau(i)}(X_{2})+d^{2}_{\tau(i)})$ (A.2)
@@ Appendix A Pseudo-stationarity
Define $\phi(\cdot):=h(\cdot)-\frac{\mu}{2}\|\cdot\|_{F}^{2}$. Then we see that $\phi$ is convex and $\nabla\phi$ is Lipschitz continuous with modulus $L-\mu$. We can now invoke [28, Theorem 2.1.5, (2.1.11)] on $\phi$ to deduce that
$0\leq(L-\mu)\langle\nabla\phi(X_{1})-\nabla\phi(X_{2}),X_{1}-X_{2}\rangle-\|\nabla\phi(X_{1})-\nabla\phi(X_{2})\|_{F}^{2}.$ (A.3)
For the right hand side of (A.3), a direct computation shows that
$(L-\mu)\langle\nabla\phi(X_{1})-\nabla\phi(X_{2}),X_{1}-X_{2}\rangle-\|\nabla\phi(X_{1})-\nabla\phi(X_{2})\|_{F}^{2}$ (A.4)
@@
We now rewrite $S_{1}$ and $S_{2}$. We start by noting that for $S_{2}$, its two summands can be rewritten as follows using A.1: for $i=1$, $2$,
$-\langle\mu X_{i}-\nabla h(X_{i}),LX_{i}-\nabla h(X_{i})\rangle=-\sum_{j=1}^{m}(\mu\sigma_{j}(X_{i})+d^{i}_{j})(L\sigma_{j}(X_{i})+d^{i}_{j}).$ (A.5)
@@
$S_{1}$ $=\langle\mu X_{1}-\nabla h(X_{1}),LX_{2}-\nabla h(X_{2})\rangle+\langle\mu X_{2}-\nabla h(X_{2}),LX_{1}-\nabla h(X_{1})\rangle$ (A.6)
=$\overset{\rm{(a)}}{=}\left\langle R_{1}\begin{bmatrix}\mu\Sigma_{1}+D_{1}&0\end{bmatrix}P_{1}^{\top},R_{2}\begin{bmatrix}L\Sigma_{2}+D_{2}&0\end{bmatrix}P_{2}^{\top}\right\rangle$
@@
where in (a) we have used A.1. Using the above display and Lemma 2.3, we see that
$S_{1}\leq\max_{\tau\in\mathfrak{P}_{m}}\sum_{i=1}^{m}(\mu\sigma_{i}(X_{1})+d^{1}_{i})(L\sigma_{\tau(i)}(X_{2})+d^{2}_{\tau(i)})+\sum_{i=1}^{m}(L\sigma_{i}(X_{1})+d^{1}_{i})(\mu\sigma_{\tau(i)}(X_{2})+d^{2}_{\tau(i)}).$
@@
Lemma A.3.
Let $\tau\in\mathfrak{P}_{m}$, $L\geq\mu\geq 0$, and define $\tau(y)=\begin{bmatrix}y_{\tau(1)}&\cdots&y_{\tau(m)}\end{bmatrix}^{\top}\in\mathbb{R}^{m}$ for all $y=\begin{bmatrix}y_{1}&\cdots&y_{m}\end{bmatrix}^{\top}\in\mathbb{R}^{m}$. Let $v^{1}$, $v^{2}$, $d^{1}$, $d^{2}\in\mathbb{R}^{m}_{+}$ with $v^{1}\neq 0$, and assume that $v^{1}$ and $v^{2}$ are sorted in descending order. Let $r=|\{i\in[m]:~{}v^{1}_{i}>0\}|$ and $r^{*}=|\{i\in[m]:~{}v^{2}_{i}>0\}|$ and assume that
$\forall i\in[r],~{}j\in[r^{*}],~{}d^{1}_{i}=\lambda,~{}d^{2}_{j}=\lambda,\quad\text{and}\quad\forall k\in[m]\setminus[r],~{}d_{k}^{1}\leq\lambda+Lv^{1}_{r}.$ (A.7)
Let $X_{1}$, $X_{2}$, $G_{1}$, $G_{2}\in\mathbb{R}^{m\times n}$ be defined as
$X_{1}=\widetilde{\rm Diag}(v^{1}),~{}X_{2}=\widetilde{\rm Diag}(\tau(v^{2})),~{}G_{1}=\widetilde{\rm Diag}(d^{1}),~{}G_{2}=\widetilde{\rm Diag}(\tau(d^{2})),$ (A.8)
and suppose that
$\sum_{i=1}^{m}(L\sigma_{i}(X_{1})+d^{1}_{i})(\mu\sigma_{\tau(i)}(X_{2})+d^{2}_{\tau(i)})+\sum_{i=1}^{m}(\mu\sigma_{i}(X_{1})+d^{1}_{i})(L\sigma_{\tau(i)}(X_{2})+d^{2}_{\tau(i)})$ (A.9)
$-\sum_{i=1}^{m}(\mu\sigma_{i}(X_{1})+d^{1}_{i})(L\sigma_{i}(X_{1})+d^{1}_{i})-\sum_{i=1}^{m}(\mu\sigma_{i}(X_{2})+d^{2}_{i})(L\sigma_{i}(X_{2})+d^{2}_{i})\geq 0$
If $G_{1}+\mu X_{1}\neq G_{2}+\mu X_{2}$, we define a quadratic function $h$ as follows:
$h(X)=$ $\frac{L}{2}\sum_{i=1}^{m}\sum_{j\neq i}^{n}X_{ij}^{2}+\frac{\mu}{2}\sum_{i=1}^{m}(X_{ii}-(X_{1})_{ii})^{2}-\langle G_{1},X\rangle$ (A.10)
$+\frac{(\langle X-X_{1},-G_{2}-\mu X_{2}+G_{1}+\mu X_{1}\rangle)^{2}}{2\langle X_{2}-X_{1},-G_{2}-\mu X_{2}+G_{1}+\mu X_{1}\rangle},$
Otherwise, we set
$h(X)=$ $\frac{L}{2}\sum_{i=1}^{m}\sum_{j\neq i}^{n}X_{ij}^{2}+\frac{\mu}{2}\sum_{i=1}^{m}(X_{ii}-(X_{1})_{ii})^{2}-\langle G_{1},X\rangle.$ (A.11)
Then $h$ is well defined, $h(\cdot)-\frac{\mu}{2}\|\cdot\|_{F}^{2}$ is convex, $\nabla h$ is Lipschitz continuous with modulus $L$, $\nabla h(X_{1})=-G_{1}$, and $\nabla h(X_{2})=-G_{2}$. Moreover, if we define $F_{r}$ as in 1.5 with the above $h$ and define $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ as
$\bar{U}=\widetilde{\rm Diag}(\sqrt{\sigma_{1}(X_{1})},\dots,\sqrt{\sigma_{r}(X_{1})})\ \ {and}\ \ \bar{V}=\widetilde{\rm Diag}(\sqrt{\sigma_{1}(X_{1})},\dots,\sqrt{\sigma_{r}(X_{1})}),$ (A.12)
then $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$.
Remark A.4.
Notice that the SVDs of $X_{1}$ and $X_{2}$ are given by
$X_{1}=I_{m}\widetilde{\rm Diag}(\sigma(X_{1}))I_{n}^{\top}\ \ {and}\ \ X_{2}=W\widetilde{\rm Diag}(\sigma(X_{2}))\begin{bmatrix}W&0\\ 0&I_{n-m}\end{bmatrix}^{\top},$
where $W\in\mathcal{P}_{m}$ corresponds to the permutation $\tau\in\mathfrak{P}_{m}$. Based on A.4, A.5 and A.6, one can show that the inequality (A.9) is equivalent to
$(L-\mu)\langle(G_{2}+\mu X_{2})-(G_{1}+\mu X_{1}),X_{1}-X_{2}\rangle\geq\|(G_{1}+\mu X_{1})-(G_{2}+\mu X_{2})\|_{F}^{2}.$ (A.13)

### v2

@@ 3 First- and second-order stationary points of $F_{r}$
Define $\phi(\cdot):=h(\cdot)-\frac{\mu}{2}\|\cdot\|_{F}^{2}$, and let us rewrite 1.8 as
$0\leq(L-\mu)\langle\nabla\phi(X_{1})-\nabla\phi(X_{2}),X_{1}-X_{2}\rangle-\|\nabla\phi(X_{1})-\nabla\phi(X_{2})\|_{F}^{2}.$ (3.8)
For the right hand side of (3.8), a direct computation shows that
$(L-\mu)\langle\nabla\phi(X_{1})-\nabla\phi(X_{2}),X_{1}-X_{2}\rangle-\|\nabla\phi(X_{1})-\nabla\phi(X_{2})\|_{F}^{2}$ (3.9)
@@
We now rewrite $S_{1}$ and $S_{2}$. We start by noting that for $S_{2}$, its two summands can be rewritten as follows using 3.6: for $i=1,2$,
$\langle\mu X_{i}-\nabla h(X_{i}),LX_{i}-\nabla h(X_{i})\rangle=-\sum_{j=1}^{m}(\mu\sigma_{j}(X_{i})+d^{i}_{j})(L\sigma_{j}(X_{i})+d^{i}_{j}).$ (3.10)
@@
$S_{1}$ $=\langle\mu X_{1}-\nabla h(X_{1}),LX_{2}-\nabla h(X_{2})\rangle+\langle\mu X_{2}-\nabla h(X_{2}),LX_{1}-\nabla h(X_{1})\rangle$ (3.11)
$\overset{\rm{(a)}}{=}\left\langle R_{1}\begin{bmatrix}\mu\Sigma_{1}+D_{1}&0\end{bmatrix}P_{1}^{\top},R_{2}\begin{bmatrix}L\Sigma_{2}+D_{2}&0\end{bmatrix}P_{2}^{\top}\right\rangle$
@@
where in (a) we have used 3.6. Using the above display and Lemma 2.3, we see that
[…]
@@
We already know that when $(\bar{U},\bar{V})$ is a stationary point of $F_{r}$ in 1.2, then $\bar{X}=\bar{U}\bar{V}^{\top}$ is a pseudo stationary point of $f$ in 1.1. The next step is to identify more structural information of $\bar{X}$ and $\nabla h(\bar{X})$ when $(\bar{U},\bar{V})$ is a second-order stationary point of $F_{r}$. The first task is to rewrite the second-order optimality conditions such that it is aligned with the SVD of $\bar{X}$. To be more precise, the second-order optimality condition can be written as $\nabla^{2}F_{r}(\bar{U},\bar{V})[U,V]^{2}\geq 0$ for all $(U,V)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$. To be more aligned with the SVDs of $(\bar{U},\bar{V})$ given in Proposition 3.2, we may rewrite it as
$\forall(U,V)\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r},~~~~\nabla^{2}F_{r}(\bar{U},\bar{V})[(RUQ^{\top},PVQ^{\top}),(RUQ^{\top},PVQ^{\top})]\geq 0,$ (3.12)
where $(R,Q)\in\mathcal{O}_{U}$ and $(P,Q)\in\mathcal{O}_{V}$. Note that using [19, Equation (3.14)], [^4] Note that [19] assumed $h$ is convex, but the derivation of Equation (3.14) there does not rely on the convexity assumption. we have
$\,\nabla^{2}F_{r}(\bar{U},\bar{V})[(RUQ^{\top},PVQ^{\top}),(RUQ^{\top},PVQ^{\top})]$ (3.13)
$=$  $2\langle R^{\top}\nabla h(\bar{X})P,UV^{\top}\rangle+\lambda(\|U\|_{F}^{2}+\|V\|^{2}_{F})+\nabla^{2}h(\bar{X})[R\left(R^{\top}\bar{U}QV^{\top}+UQ^{\top}\bar{V}^{\top}P\right)P^{\top}]^{2}.$
Next, we introduce a natural partition for matrices of sizes $\mathbb{R}^{m\times r}$ and $\mathbb{R}^{n\times r}$. Let $(\bar{U},\bar{V})\in\mathbb{R}^{m\times r}\times\mathbb{R}^{n\times r}$ satisfy $\mathrm{rank}(\bar{U})=\mathrm{rank}(\bar{V})$ (which holds in particular when $(\bar{U},\bar{V})$ is a stationary point of $F_{r}$, thanks to Proposition 3.2). Denote this common rank by $s=\mathrm{rank}(\bar{U})=\mathrm{rank}(\bar{V})\leq r$. We can then partition any matrices $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{n\times r}$ into the following block form:
where $U_{11},V_{11}\in\mathbb{R}^{s\times s}$, $U_{12},V_{12}\in\mathbb{R}^{s\times(r-s)}$, $U_{21}\in\mathbb{R}^{(m-s)\times s}$, $V_{21}\in\mathbb{R}^{(n-s)\times s}$, $U_{22}\in\mathbb{R}^{(m-s)\times(r-s)}$, $V_{22}\in\mathbb{R}^{(n-s)\times(r-s)}$. Note that when $\bar{U}$ and $\bar{V}$ are of full rank, i.e., $s=r$, the matrices $U_{12}$, $U_{22}$, $V_{12}$ and $V_{22}$ are null. Now, we are ready to state the following characterization of second-order stationarity of $F_{r}$ in 1.2.
Proposition 3.6 (Second-order stationarity).
  1. (i)
  2. (ii)
    For any $U\in\mathbb{R}^{m\times r}$ and $V\in\mathbb{R}^{n\times r}$, it holds that [^5] Here, we use the partition 3.14 with respect to $(\bar{U},\bar{V})$; this is well defined because $\sigma(\bar{U})=\sigma(\bar{V})$ holds in (i).
    where $\Sigma=\mathrm{Diag}(\sigma_{1}(\bar{U}),\dots,\sigma_{s}(\bar{U}))\in\mathbb{R}^{s\times s}$, and $D=\widetilde{\rm Diag}(d_{s+1},\dots,d_{m})\in\mathbb{R}^{(m-s)\times(n-s)}$ with $d_{i}$ given in (i).
Moreover, if $s=\mathrm{rank}(\bar{U})<r$, then (i) and (ii) imply that [^6] We would like to point out this last claim (i.e., “if $s=\mathrm{rank}(\bar{U})<r$, then (i) and (ii) imply that $\|\nabla h(\bar{U}\bar{V}^{\top})\|_{2}\leq\lambda$.”) has been proved in [3, Theorem 1]. We include its proof for completeness. $\|\nabla h(\bar{U}\bar{V}^{\top})\|_{2}\leq\lambda$.